IDENTIFICATION OF THE [12.3] ^{2}Σ^{+} − B^{2}Σ^{+} TRANSITION OF NiCl

\underline{LEAH O’BRIEN} and COREY RICE, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared system of NiCl at 10 496 cm^{−1}has been recorded and identified as the [12.3] ^{2}Σ^{+} − B^{2}Σ^{+} transition of NiCl. Improved molecular constants will be presented. Unusual intensity patterns are observed in the spectrum indicating a large perpendicular component of the transition dipole moment.